OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!HELP PLEASE!!!!! What does Solve- 5x=30 mean?

OpenStudy (diyadiya):

Solve for x or Find the value of x

OpenStudy (diyadiya):

so what do you do here?

OpenStudy (diyadiya):

you can isolate x by dividing both sides by 5

OpenStudy (anonymous):

So its x = 6?

OpenStudy (diyadiya):

YES!!

OpenStudy (anonymous):

5x/5=30/5 So X=6

OpenStudy (anonymous):

YAY im smart!!!! ahhahaha :D

OpenStudy (diyadiya):

lol yes you are xD

OpenStudy (anonymous):

y over 3 = 9

OpenStudy (anonymous):

wat do i do for that hahah

OpenStudy (anonymous):

@Diyadiya

OpenStudy (diyadiya):

\[\Large \frac{y}{3}=9\]Multiply both sides by 3 \[\Large \frac{y}{3}\times 3 = 9 \times 3 \]\[\Large \frac{y}{\cancel{3}}\times \cancel{3} = 9 \times 3 \] SO y = ?

OpenStudy (anonymous):

27?.......

OpenStudy (diyadiya):

YES!! :) btw did you understand ?

OpenStudy (anonymous):

Yeah i did hahah thanx

OpenStudy (anonymous):

lol how old are you?

OpenStudy (diyadiya):

Glad to help :) i'm Fine Thanks :)

OpenStudy (diyadiya):

oh old

OpenStudy (diyadiya):

17

OpenStudy (anonymous):

hahah

OpenStudy (anonymous):

lol im 14

OpenStudy (diyadiya):

lol ok

OpenStudy (diyadiya):

* btw these questions should be in the chat :D *